OpenStudy (anonymous):

(2x+y)^2 How do you solve this?

OpenStudy (jdoe0001):

\(\bf (2x+y)^2\implies (2x+y)(2x+y)\) FOIL

OpenStudy (anonymous):

\[(2x+y)^{2} = 4x ^{2}+y ^{2} + 4xy\]

OpenStudy (anonymous):

\[(a+b)^{2}= a ^{2}+b ^{2}+2ab\]

OpenStudy (texaschic101):

(2x + y)^2 = (2x + y)(2x + y) = 4x^2 + 2xy + 2xy + y^2 (combine like terms) 4x^2 + 4xy + y^2